OpenStudy (anonymous):

Solve, using the substitution method. y = 3x + 5 4x – y = 5 A. (10, 35) B. (15, 10) C. There are an infinite number of solutions. D. There is no solution.

OpenStudy (the_fizicx99):

4x –(3x + 5) = 5 4x -3x -5 = 5 1x - 5 = 5 +5 +5 x = 10 4(10) + y = 5 40 + y = 5 -40 -40 y = 35 (10,35)

OpenStudy (anonymous):

thanks a bunch

OpenStudy (the_fizicx99):

yw

OpenStudy (anonymous):

1 more please

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

Solve, using the substitution method. y + 2x = 7 14 – 4x = 2y A. The solution is (1, 5) B. The solution is (21, 0) C. There are an infinite number of solutions. D. There is no solution.

OpenStudy (anonymous):

aha have fun fizi

OpenStudy (the_fizicx99):

C

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

Here ya go :)

OpenStudy (anonymous):

the answer is A

OpenStudy (anonymous):

thanks @lovethecreation